OpenStudy (anonymous):

What are the solutions of x2 + 8x − 20 = 0

OpenStudy (anonymous):

This is a quadratic equation....there are several ways you can solve for x but note that you will have two values for x.

OpenStudy (anonymous):

I think in this case we can complete the square

OpenStudy (anonymous):

(x+a)(x+b) a and b are two number that if you multiply them together you get -20 but if you add them you will get 8

OpenStudy (anonymous):

(x+a)(x+b)=0 sorry I forgot to set it to zero

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

or it you just understand this at all you can always use the quadratic formula. \[\begin{array}{*{20}c} {x = \frac{{ - b \pm \sqrt {b^2 - 4ac} }}{{2a}}} & {{\rm{when}}} & {ax^2 + bx + c = 0} \\ \end{array} \] \[x^2 + 8x − 20 = 0\] in this case a=1 b=8 c=-20